Solve the equation logxe2e = eInx-e


1
Expert's answer
2021-08-31T11:54:26-0400

"\log _{x} e^{2 e}=e \ln x-e
\\\Rightarrow \log _{x} e^{2 e}=e \log _{e} x-e
\\\Rightarrow \log _{x} e^{2 e}=\frac{e}{\log _{x} e}-e
\\\Rightarrow 2 \mathrm{e} \cdot \log _{x} e=e\left(\frac{1}{\log _{x} e}-1\right)
\\\Rightarrow 2 \log _{x} e=\left(\frac{1}{\log _{x} e}-1\right)
\\Let \log _{x} e=t
\\So,
\\2 t=\frac{1}{t}-1 \\\Rightarrow 2 t^{2}+t-1=0 \\\Rightarrow 2 t^{2}+2 t-t-1=0
\\\Rightarrow 2 t(t+1)-1(t+1)=0 \\\Rightarrow(2 t-1)(t+1)=0 \\\Rightarrow 2 t-1=0\ or\ t+1=0 \\\Rightarrow t=\frac{1}{2},-1
\\As \log _{x} e=t"

We know that,

"{ \log _{x} e>0}"

"So, \log _{x} e \neq-1
\\Now, \log _{x} e=\frac{1}{2} \\\Rightarrow e=x^{\frac{1}{2}}\\\Rightarrow x=e^2
\\Ans: x=e^2"